OpenStudy (anonymous):

CAN SOME ONE HELP ME WITH MY 06.05 FREQUENCY TABLE QUIZ?

undeadknight26 (undeadknight26):

i took the same test.

undeadknight26 (undeadknight26):

What do u need help with?

OpenStudy (anonymous):

Can u help i dont understand

OpenStudy (anonymous):

everything

undeadknight26 (undeadknight26):

Well...i can help u understand something you got wrong maybe...but im not really a teacher...

OpenStudy (anonymous):

ok that will work

undeadknight26 (undeadknight26):

Can you maybe copy a question please? Thanks :D

OpenStudy (anonymous):

Families were surveyed to see how they spent summer vacation. The results are as follows: One or more children below age 16—Camping: 30; Traveling out of state: 16; Activities in local area: 44 All children 16 or older—Camping: 22; Traveling out of state: 60; Activities in local area: 28 Which of the following is a correct two-way relative frequency table for the data? is it what i think it is? like just put the numbers in the table

undeadknight26 (undeadknight26):

Can you just save the graphs to ur computer and upload it? That would be great :D

OpenStudy (anonymous):

um ill try my com is stupid.

undeadknight26 (undeadknight26):

mine is too...can't even game right...

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

thx

undeadknight26 (undeadknight26):

i can't pull it up :O

OpenStudy (anonymous):

ugh! hold on....

OpenStudy (anonymous):

High school students from grades 9-10 and 11-12 were asked to choose the kind of band to have play at a school dance: Rap, Rock, or Country. Their choices were as follows: Grades 9-10—Rap: 40; Rock: 30; Country: 55 Grades 11-12—Rap: 60; Rock: 25; Country: 35 Which of the following is a correct two-way frequency table for the data? Band preference for school dance Rap Rock Country Row totals Grades 9-10 0.28 0.21 0.38 0.86 Grades 11-12 0.41 0.17 0.24 0.83 Column totals 0.69 0.38 0.62 1 Band preference for school dance Rap Rock Country Row totals Grades 9-10 0.40 0.55 0.61 0.86 Grades 11-12 0.60 0.45 0.39 0.83 Column totals 1 1 1 1 Band preference for school dance Rap Rock Country Row totals Grades 9-10 40 30 55 125 Grades 11-12 60 25 35 120 Column totals 100 55 90 245 Band preference for school dance Rap Rock Country Row totals Grades 9-10 0.32 0.24 0.44 1 Grades 11-12 0.50 0.21 0.29 1 Column totals 0.69 0.38 0.62 1

undeadknight26 (undeadknight26):

lets see... 40 + 30 + 55 = x 40/x = a

undeadknight26 (undeadknight26):

did you do that?

OpenStudy (anonymous):

oh srry mom called me wut?

OpenStudy (anonymous):

wich table is it?

undeadknight26 (undeadknight26):

well im trying to teach you how to do it...im honestly not very good at these either o.o

OpenStudy (anonymous):

ok srry :(

OpenStudy (anonymous):

uum can u try im patent

undeadknight26 (undeadknight26):

ok so: 40 + 35 + 55 = 130 40/130 = .31 35/130 = .27 55/130 = .42

undeadknight26 (undeadknight26):

Oh wait im doing a relative o.o its C because they just wanted the graph i believe...

OpenStudy (anonymous):

ok thx can u help with another ;/

undeadknight26 (undeadknight26):

I really look over the easy questions...and sure i can...not 100% accurate with my answers though...

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

Using the following table, what term best describes the entry of 62%? Which location would you most like to visit? Aspen, Colorado New York, New York Row totals Male students 22% 28% 50% Female students 16% 34% 50% Column totals 38% 62% 100% Joint frequency Joint relative frequency Marginal frequency Marginal relative frequency

undeadknight26 (undeadknight26):

That would be a marginal frequency...

OpenStudy (anonymous):

thx another?

undeadknight26 (undeadknight26):

surez.

OpenStudy (anonymous):

The table shows the movie preferences of high school and college students. Movie Preference Action Drama Comedy Row totals High School 40 23 12 75 College 20 51 28 99 Column totals 60 74 40 174 Julie mistakenly calculated the conditional relative frequency for college students who like comedy as being 70%. What did Julie actually calculate and what is the correct answer? She calculated the joint relative frequency for college/comedy. The correct value of conditional relative frequency for college students who like comedy is 28%. She calculated the conditional relative frequency for comedy films that are liked by college students. The correct value of conditional relative frequency for college students who like comedy is16%. She calculated the conditional relative frequency for comedy films that are liked by college students. The correct answer is 28%. She calculated the joint relative frequency for college/comedy. The correct answer is 16%.

OpenStudy (anonymous):

The table shows the movie preferences of high school and college students. Movie Preference Action Drama Comedy Row totals High School 40 23 12 75 College 20 51 28 99 Column totals 60 74 40 174 Julie mistakenly calculated the conditional relative frequency for college students who like comedy as being 70%. What did Julie actually calculate and what is the correct answer? She calculated the joint relative frequency for college/comedy. The correct value of conditional relative frequency for college students who like comedy is 28%. She calculated the conditional relative frequency for comedy films that are liked by college students. The correct value of conditional relative frequency for college students who like comedy is16%. She calculated the conditional relative frequency for comedy films that are liked by college students. The correct answer is 28%. She calculated the joint relative frequency for college/comedy. The correct answer is 16%.

undeadknight26 (undeadknight26):

well its either A or B...thats a hard one to me...

OpenStudy (anonymous):

same enie......meanie.... mineie .....moe

undeadknight26 (undeadknight26):

lol XD

OpenStudy (anonymous):

ugh my profile pic wont upload!!! (u mad bro) YES!!!

undeadknight26 (undeadknight26):

lol XD you went to settings and clicked change pic?

OpenStudy (anonymous):

yes it says........uploading...uploading....up.AHH!!

undeadknight26 (undeadknight26):

lol how big is the file??? takes me like 10 secs to upload mine....maybe you should try doing it again with a smaller file?

OpenStudy (anonymous):

ugh my profile pic wont upload!!! (u mad bro) YES!!!

OpenStudy (anonymous):

ugh i hate the glitches on this thing give me five mins? got to reboot.

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

@Tateofgrace u talking in this seriously pisses me of

undeadknight26 (undeadknight26):

Calm down there mate XD

OpenStudy (anonymous):

@undeadknight26 could you help me with some questions on this same test? it would mean so much omg

OpenStudy (anonymous):

or the 04.01 test??

undeadknight26 (undeadknight26):

@memex3lol Yeah sure! Im really rusty on this though!

OpenStudy (anonymous):

Which statement is true about the sum square root of \[\sqrt{50}+\sqrt{18}\] A. It is rational and equal to 2. B. It is rational and equal to 8. C. It is irrational and equal to \[2 \sqrt{2}\] D. It is irrational and equal to \[8 \sqrt{2}\]